A company is considering two capital expenditure proposals. Both proposals are for similar products andboth are expected to operate for four years. Only one proposal can be accepted.The following information is available.Profit/(loss)

Proposal A Proposal B

Initial investment $45,000 $ 46,000

Year 1 6,500 4,500

Year 2 3,500 2,500 Year 3 13,500 4,500 Year 4 Loss (1,500) profit14,500 Estimated scrap value

at the year end 4 4,000 4,000 Depreciation is charged on the straight line basis

Calculate the following for both proposals

I) The payback period to one decimal place

ii) The average rate of return on initial investment to the decimal place



1
Expert's answer
2021-07-15T10:39:31-0400

Solution:

i). The payback period:

First, calculate depreciation and add to the profit or loss for the periods to determine the net cash flows:

Depreciation = "\frac{Cost - Scrap\; value}{No. \; of\; years} = \frac{46,000 - 4,000}{4} = 10,500"


The payback period = "Years\; before\; full \;recovery + \frac{Unrecovered \;cost\;at\;the\;start\;of\;the\;year}{Cash\; flow \;during\; the\; year}"


The payback period for Project A and Project B calculations is as follows:

 

 


Payback period project A:

The payback period for Project A = 2.6, that is, 2 years and 6 months.

 

Payback period project B:

The payback period for Project B = 3.1, that is, 3 years and 1 month.



ii.). The average rate of return on initial investment:

Project A ARR:

The average rate of return (ARR) = "\frac{Annual\; average\;profits}{(Initial\;investment - scrap \;value)}\times 100\%"


Annual average profits = "\frac{6500 + 3500 + 13500 - 1500}{4} = 5,500"


Initial investment – scrap value = 46,000 – 4000 = 42,000


The average rate of return (ARR) = "\frac{5,500}{42,000} \times 100\% = 13.1\%"


The average rate of return (ARR) for Project A = "13.1\%"


Project B ARR:

The average rate of return (ARR) = "\frac{Annual\; average\;profits}{(Initial\;investment - scrap \;value)}\times 100\%"


Annual average profits = "\frac{4500 + 2500 + 4500 + 14500}{4} = 6,500"


Initial investment – scrap value = 46,000 – 4000 = 42,000


The average rate of return (ARR) = "\frac{6,500}{42,000} \times 100\% = 15.5\%"


The average rate of return (ARR) for Project B = "15.5\%"